OpenStudy (anonymous):

Use the three steps to solve the problem. One of two supplementary angles is five times the other. Find the angles. (Recall that supplementary angles are angles whose sum is equal to 180 .)

OpenStudy (ajprincess):

Let one of the angle be x. Then the other angle will be 5x. Since the sum of supplementary angles is 180 x+5x=180 6x=180 Solve for x. Can u do it @KaiteDID?

OpenStudy (anonymous):

174?

OpenStudy (ajprincess):

hmm no. 6x=180 6x/6=180/6 x=30. So the value of one of the angles is 30. Can u find the other angle?

OpenStudy (anonymous):

150?

OpenStudy (ajprincess):

yup u r right:)

OpenStudy (anonymous):

thank you for u help

OpenStudy (ajprincess):

yw:)